OpenStudy (anonymous):

Just a simple chemistry question (I am posting it here as the Chemistry study group is DEAD)

OpenStudy (anonymous):

Calculate the concentration of water and table salt, when 250 grams of table salt NaCl is in 3 litres of liquid. Table salt has a molecular mass of 58,5 g/mol. Give the answer in two decimal precision.

OpenStudy (anonymous):

I know the formula for moles and concentration, which is mol=mass/RMM and conc=mol/volume I did the calculations, but I am still not sure if I am correct, I got 1.42 as the final answer. And I don't know what units it should be in, is it Molarity (that would be 1.42M) or mol/l, or mol/dm^3?? so, if anyone can tell me if I am right about the answer being 1.42 and also what units I should use, it would be great.

OpenStudy (anonymous):

Oh, are you looking for molars?

OpenStudy (anonymous):

mol per liter?

OpenStudy (anonymous):

\[\frac{250g}{58.5g/mol}=4.27mol\]so the concentration will be\[\frac{4.27mol}{3L}=1.42M\]

OpenStudy (anonymous):

great, so I was right :) so I always have to use Molars right? M?

OpenStudy (anonymous):

Depends on what they want. Typically it's that. Another kind of concentration is the 'molal' which is moles of solute ----------------- mass of solvent in kg

OpenStudy (anonymous):

ok, btw, do you know what exactly is constant forward carrying force? is it just mass*acceleraltion? or something involving momentum?

OpenStudy (anonymous):

I've never heard Newton's Second Law of Motion called 'constant forward carrying force'...

OpenStudy (anonymous):

\[F=\frac{dp}{dt}\]where p is momentum and F is force. This is the 'proper' definition of force. If then, you take momentum, p, as\[p=mv\]then\[F=\frac{dp}{dt}=\frac{d(mv)}{dt}=m \frac{dv}{dt}=m a\]

OpenStudy (anonymous):

oh, ok, because I have this simple question and I just wanted to be sure..

OpenStudy (anonymous):

When there's no speeds are small compared to speed of light.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Motorcyclist (total mass 320 kg) stops from a velocity of 36 km/h in 25 s without breaking. The same motorcyclist accelerates in the same conditions from full stop to velocity of 126 km/h in 6 s. Calculate the constant forward carrying force of the engine, if the forces restricting the motion remain constant. Give the answer in 1 N precision. so, I used the newton's second law of motion and found the answer to be 1867N. Is that right? my calculations were like this: converted 126km/h to 35m/s F=\frac{dp}{dt} F=\frac{(35*320)}{(6)} F=1866.66 F=1867N am I right? or did I miss something important?

OpenStudy (anonymous):

gtg, will be back in like 15 min

OpenStudy (anonymous):

Hey, because I'm fried (it's 230am here) and didn't want to leave you hanging, I found this website that has the question you just asked AND solution: http://www.algebra.com/algebra/homework/testmodule/testmodule.faq Go nuts!

OpenStudy (anonymous):

I have to sign out - my eyes are burning -.- Study hard!

OpenStudy (anonymous):

ok, thanks !!! You are a real HERO!